OpenStudy (anonymous):

Inverse Laplace transformation of 1/(s^2-s) could someone show me how to do this

OpenStudy (phi):

I would factor the denominator to get s(s-1) then use partial fraction decomposition to find A and B so that \[ \frac{1}{s(s-1)}= \frac{A}{s} + \frac{B}{s-1} \]

OpenStudy (anonymous):

lol. I was thinking of manipulating the function. thanks